# Dynamics of the sine-Gordon equation on tadpole graphs

### v1

@@ 4.1 A priori restrictions on the parameter $Z$
$-Z>\frac{\phi^{\prime}_{2}(L)}{\phi_{2}(L)}=-\frac{1}{c_{2}}\frac{y}{(1+y^{2})\arctan(y)}\equiv f(y),\;\;\quad y>0.$ (4.2)
@@
From (4.2) we get initially that $Z<\frac{1}{c_{2}}$ because $f(y)>-\frac{1}{c_{2}}$ for all $y>0$. Now, for $a\leqq 0$ we have $y=e^{-a/c_{2}}\geqq 1$. Then $-Z>f(y)\geqq f(1)=-\frac{2}{\pi c_{2}}$, and so $Z<\frac{2}{\pi c_{2}}$. For $a>0$, we have $y<1$ and so $f(y)<f(1)$. Thus, for $a\to 0^{+}$, $f(y)\to f(1)$, and therefore $f(1)\leqq-Z$. But, for $a=0$, from (4.2) follows $-Z>f(1)$. In conclusion, we need to have the restriction $Z\in(-\infty,\frac{2}{\pi c_{2}})$. ∎
@@ 4.2 Existence of positive single-lobe states (subluminal librations)
$-c_{1}^{2}\phi^{\prime\prime}_{1}(x)+\sin(\phi_{1}(x))=0,\qquad x\in[-L,L],$ (4.3)
with a single-lobe on $[-L,L]$, namely, $\phi_{1}$ is even, the maximum of $\phi_{1}$ is achieved at a single internal point of $[-L,L]$, namely, $x=0$, and $\phi^{\prime}(x)<0$ on $(0,L]$ (see Figure 3 and the phase-plane in Figure 5).
Figure 5: Phase-plane of a positive single-lobe kink state for (1.6). Here we take $c_{1}=c_{2}=1$ and the translation variable $\xi=x-ct$ with $c=0$ (subluminal waves). The picture shows the phase plane of the solutions $\phi=\phi(\xi)$ for (1.6). The separatrices are represented in red solid lines and the librations are the periodic waves inside the separatrices (in light gray). The single-lobe kink profile (depicted with blue thick arrow lines) is composed of a periodic libration, symmetric in $[-L,L]$ and connecting $\phi=\pi$ with itself, glued with a decaying kink connecting $\phi=\pi$ to $\phi=0$ along a separatrix (color online).
@@
$-\frac{c_{1}^{2}}{2}[\phi^{\prime}_{1}(x)]^{2}-\cos(\phi_{1}(x))=-\cos(\phi_{1}(0))\Longrightarrow-\frac{c_{1}^{2}}{2}[\phi^{\prime}_{1}(x)]^{2}+1-\cos(\phi_{1}(x))=E,\;\;x\in[-L,L],$ (4.4)
@@
In the sequel we obtain two explicit formulae for the profile $\phi_{1}$ in (4.6) being an even single-lobe profile on $[-L,L]$ such that $\phi_{1}^{\prime}(x)<0$ for $x\in(0,L)$ and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. As we shall see, these formulae will depend on the position of $\phi_{2,a}(L)$ with regard to $\pi$, where the shift-parameter $a$ determines the kink-soliton profile in (1.8) (see Figure 3).
@@
Let $L$ and $c_{1}$ be fixed positive constants. We consider a priori $a<0$ in (1.8) which determines the kink-soliton profile $\phi_{2,a}$ (therefore $\phi_{2,a}(L)>\pi$). Suppose that $\phi_{1}$ satisfies equation in (4.3) and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. Then, there is a family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (4.3) of subluminal type with a single-lobe profile on $[-L,L]$ given by,
@@
Next, from condition $\phi_{1}(x)>\pi$ on $[-L,L]$, the property $\operatorname{sn}(2K(k);k)=0$ and (4.6), we need to have $\frac{x}{c_{1}}+K<2K$ for all $x\in[0,L]$ (see Figure 4 above). Therefore, a priori we get that modulus $k$ satisfies $K(k)>\frac{L}{c_{1}}$. Next, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$ we obtain immediate a family of single-lobe solutions $k\to\phi_{1,k}$ for any $k\in(0,1)$. For $\frac{L}{c_{1}}>\frac{\pi}{2}$, we consider $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$, then we have the family of single-lobe profiles $k\in(k_{0},1)\to\phi_{1,k}$ defined in (4.7). Lastly, from (4.3) we get immediately $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$. ∎
@@
We consider $L=\pi$, $c_{1}=1$, then $K(k)=\pi$ for $k\approx 0.9843\equiv k_{0}$, and so for $k>k_{0}$ we get the family of subluminal solutions $k\to\phi_{1,k}$ in (4.7) with a single-lobe profile on $[-L,L]$. Figure 4 shows the profile of $\phi_{1,k}$ for $k^{2}=0.98$.
@@
Let $L,c_{1}$ be fixed positive constants. We consider a priori $a>0$ in (1.8) which determines the kink-soliton profile $\phi_{2,a}$ (with $\phi_{2,a}(L)<\pi$). Suppose that $\phi_{1}$ satisfies equation in (4.3) and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. Then, there is a family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (4.3) of subluminal type with a single-lobe profile on $[-L,L]$ given by
@@
From $a>0$, $\phi_{1}(0)>\pi$ and the boundary condition $\phi_{1}(L)=\phi_{2,a}(L)<\pi$, we have a unique $b\in(0,L)$, $b=b(k)$, such that $\phi_{1}(b)=\pi$, and so for $x\in[0,b]$, $\phi_{1}(x)\in[\pi,\phi_{1}(0)]$, and for $x\in[b,L]$, $\phi_{1}(x)\in[\phi_{1}(L),\pi]$. Thus, from (4.6) we get the single-lobe profile in (4.8). Next, from $\phi_{1,k}(b)=\pi$ and the strictly-decreasing property of $\phi_{1,k}$ on $[0,L]$ we need to have $b(k)=c_{1}K(k)$ (see Figure 7 above). Therefore, $K(k)=\frac{b}{c_{1}}<\frac{L}{c_{1}}$. Moreover, from (4.3) we obtain $\phi^{\prime\prime}_{1}(b)=0$, $\phi^{\prime\prime}_{1}(x)<0$ for $x\in[0,b)$ and $\phi^{\prime\prime}_{1}(x)>0$ for $x\in[b,L]$ (see Figure 9 below). ∎
@@
We consider $L=\pi$, $c_{1}=1$, then $K(k)=\pi$ for $k\approx 0.984432\equiv k_{0}$, and so for $k\in(0,k_{0})$ we get the family of subluminal librational solutions $k\to\phi_{1,k}$ in (4.8) with a single-lobe profile on $[-L,L]$ with $b=K(k)$. Figures 8 and 8 show the parts of $\phi_{1,k}$ in (4.8) when they are considered on the complete interval $[0,\pi]$ with $k=0.5$. Figure 9 shows the full profile of $\phi_{1,k}$ in (4.8) with $k=0.5$, $b=K(0.5)$ and it which is based on gluing the two components for $\phi_{1,k}$ in (4.8) from the profiles in Figures 8 and 8. We note that the “corners” (namely, the points where the derivative does not exist) in Figures 8 and 8 occur simultaneously at $x=b=K(0.5)$.
@@
Figure 9: Graph of the single-lobe solution (in solid blue line), composed by the extension of $\phi_{1,k}$ in (4.8) with $c_{1}=1$, $k=0.5$ to the whole interval $x\in[-\pi,\pi]$. The glueing of both solutions occurs at $x=b$ and $x=-b$ with $b\approx 1.85407$ (color online).
@@ 4.3 Existence of a single-lobe kink state
Let $L,c_{1}$ fixed positive constants. For $a<0$ we consider the kink-soliton profile $\phi_{2,a}$ in (1.8) with $\phi_{2,a}(L)>\pi$ and the single-lobe profile $\phi_{1,k}$ in (4.7). We have a single-lobe kink state, namely $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, in the following two cases:
@@
The proof of Proposition 4.7 below shows formulae for $m_{0},m_{1},m_{2}$. Since $\tanh(K(k))>k$ for all $k\in(0,1)$, then in the case $(1)(i)$ above we necessarily have $\frac{c_{1}}{2c_{2}}-k_{0}>0$.
Proof of Proposition 4.7.
Let $a<0$. The idea in the following is to glue the subluminal-profile $\phi_{1,k}$ in (4.7) with $\phi=\phi_{2,a}$ at the vertex $\nu=L$ such that $\phi_{1,k}(L)=\phi_{2,a}(L)$ and with $k$ satisfying a priori $K(k)>\frac{L}{c_{1}}$.
Indeed, by formulae (1.8) and (4.7) we obtain the following first formula for the shift parameter $a=a(k)$,
@@
where $\mathrm{dn}(\cdot;k)$ and $\mathrm{cn}(\cdot;k)$ represent the Jacobian elliptic functions of dnoidal and cnoidal type, respectively (see Figure 7). Here, $k^{\prime}$ is defined by $k^{\prime}=\sqrt{1-k^{2}}$. In the following, we show formula in (4.9). Indeed, for $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$ we get from (4.6) that $\cos(4\beta)=-1+2k^{2}sn^{2}\Big{(}\frac{L}{c_{1}}+K;k\Big{)}$. Then, from $\cos(4\beta)=2\cos^{2}(2\beta)-1$ follows $k^{2}\operatorname{sn}^{2}(\frac{L}{c_{1}}+K;k)=\cos^{2}(2\beta)$ Thus, since $2\beta\in[\frac{\pi}{2},\pi]$ and $\frac{L}{c_{1}}+K<2K$ we get $\cos(2\beta)=-k\operatorname{sn}(\frac{L}{c_{1}}+K;k)$ (see, for instance, Figure 4). Moreover, from $\cos^{2}(2\beta)+\sin^{2}(2\beta)=1$ and $1-k^{2}\operatorname{sn}^{2}=\mathrm{dn}^{2}$ we also obtain $\sin(2\beta)=\mathrm{dn}(\frac{L}{c_{1}}+K;k)$. Therefore,
@@
and so we get the property that the shift-value mapping $k\to a(k)$ will be smooth. Formula (4.11) shows that always we can glue a librational profile with a kink-profile because $k^{\prime}\leqq\mathrm{dn}(x;k)$ for all $x$ and $k$.
In the sequel we determine the correct values of $k$ and $Z$ such that $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, namely, such that we have the relation $2\phi^{\prime}_{1}(L)=\phi^{\prime}_{2}(L)+Z\phi_{2}(L)$. We divide this study into several cases.
$\bullet$ Case (1). Let $\frac{L}{c_{1}}>\frac{\pi}{2}$. Consider the unique $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$. Then for $k>k_{0}$ it follows that $K(k)>\frac{L}{c_{1}}$. Now, from properties of the Jacobi elliptic functions, we obtain from (4.11) the following limits:
$\lim_{k\to 1}\operatorname{sech}\Big{(}\frac{a(k)}{c_{2}}\Big{)}$ $=\lim_{k\to 1}\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)}=\frac{0}{\operatorname{sech}(L/c_{1})}=0,$ (4.12)
$\frac{a(k)}{c_{2}}=-\ln\Big{(}\frac{\mathrm{dn}(L/{c_{1}};k)+k\mathrm{cn}(L/{c_{1}};k)}{k^{\prime}}\Big{)}\equiv j(k).$ (4.13)
Then, numerical simulations always show that $j(k)$ in (4.13) is strictly decreasing for $k\in(k_{0},1)$ and so $a^{\prime}(k)<0$, for $k\in(k_{0},1)$ (see Figure 11 for the graph of $j=j(k)$, $k\in(k_{0},1)$ with $c_{1}=1$, $L=\pi$ and $k_{0}\approx 0.984432$). We note that by using formula (710.53) in [22] and deriving with regard to $k$ both sides of equality in (4.11) we can also see that $a(k)$ is strictly decreasing after long calculations.
Figure 11: Graph of the function $j=j(k)$ in (4.13) for $k\in(k_{0},1)$, with $c_{1}=1$, $L=\pi$ and where $k_{0}\approx 0.984432$.
$\phi^{\prime}_{1}(L)=\frac{2k}{c_{1}}\mathrm{cn}(L/c_{1}+K;k)=-\frac{2kk^{\prime}\operatorname{sn}(L/c_{1};k)}{c_{1}\mathrm{dn}(L/c_{1};k)}.$ (4.14)
From (1.8) and (4.11) we get for $\phi_{2}=\phi_{2,a}$,
$\phi^{\prime}_{2}(L)=-\frac{2}{c_{2}}\operatorname{sech}\Big{(}\frac{a}{c_{2}}\Big{)}=-\frac{2}{c_{2}}\frac{k^{\prime}}{\mathrm{dn}(L/c_{1};k)}.$ (4.15)
$\frac{k^{\prime}}{c_{1}\mathrm{dn}(L/c_{1};k)}\Big{[}\frac{c_{1}}{2c_{2}}-k\operatorname{sn}(L/c_{1};k)\Big{]}=Z\beta,$ (4.16)
$Z=\frac{\operatorname{sech}(a/c_{2})}{c_{1}\arctan(e^{-a/c_{2}})}\Big{[}\frac{c_{1}}{2c_{2}}-k\operatorname{sn}(L/c_{1};k)\Big{]}\equiv H(k).$ (4.17)
$H(k)=Z,$ (4.18)
for $Z\in(-\infty,\frac{2}{\pi c_{2}})$ fixed, and some $k\in(k_{0},1)$ with $a=a(k)$ solving (4.11). Next, we consider the following cases:
  1. (I)
    Let $Z>0$. Then, a priori, $k$ needs to satisfy $k\operatorname{sn}(L/c_{1};k)<\frac{c_{1}}{2c_{2}}$ for solving (4.18). Thus, from basic properties of the Jacobi elliptic functions we get the following limits
    $\lim_{k\to 1}k\operatorname{sn}(L/c_{1};k)$ $=\tanh(L/c_{1}),$ (4.19)
    Moreover, the mapping
    $k\in(k_{0},1)\to g(k)\equiv k\operatorname{sn}(L/c_{1};k),$ (4.20)
    is strictly increasing (we note that for $k<k_{0}$ it is possible to have $g(k)\leqq 0$ and with $g$ having some oscillations; see Figures 12 and 13), and so we get
    Figure 12: Graph of the function $g(k)$ defined on (4.20) for $k\in(0,1)$, with $c_{1}=1$, $L=2.5\pi$.
    Let us consider the following subcases:
    1. (i)
      Suppose $\frac{c_{1}}{2c_{2}}\geqq\tanh(L/c_{1})$. Then $H(k)>0$ for $k\in(k_{0},1)$. Now, from (4.12) and properties of the Jacobi elliptic functions (cf. [22]), we obtain
      $\lim_{k\to 1}H(k)$ $=\frac{0}{\arctan(\infty)}\Big{[}\frac{c_{1}}{2c_{2}}-\tanh(L/c_{1})\Big{]}=0,$ (4.21)
    2. (ii)
      Suppose $k_{0}<\frac{c_{1}}{2c_{2}}<\tanh(L/c_{1})$. Then, there is a unique $\gamma\in(k_{0},1)$ with $g(\gamma)=\frac{c_{1}}{2c_{2}}$ and such that for $k\in(k_{0},\gamma)$ follows $H(k)>0$. From $\lim_{k\to\beta}H(k)=0$ and (4.21), we have for $Z\in\big{(}0,\frac{4}{\pi c_{1}}\big{[}\frac{c_{1}}{2c_{2}}-k_{0}\big{]}\big{)}$ that there is a unique $k_{Z}\in(k_{0},\beta)$ with $H(k_{Z})=Z$. This shows $(1)(ii)$.
  2. (II)
    Now let $Z<0$. From (4.17) follows that $k$ a priori needs to satisfy $\frac{c_{1}}{2c_{2}}<k\operatorname{sn}(L/c_{1};k)<\tanh(L/c_{1})$. Let us examine the following subcases:
    1. (iii)
    2. (iv)
    3. (v)
  3. (III)
    Now suppose that $Z=0$. From (4.17) we need to have $H(k)=0$ and so $\frac{c_{1}}{2c_{2}}=k\operatorname{sn}(L/c_{1};k)$ is true if and only if $\frac{c_{1}}{2c_{2}}\in(k_{0},\tanh(L/c_{1}))$ and in this case there is a unique $k\in(k_{0},1)$ such that $H(k)=0$. This shows $(1)(vi)$.
$\bullet$ Case (2). Now consider the case $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$. Then $K(k)>\frac{L}{c_{1}}$ for all $k\in(0,1)$. Thus, since $\mathrm{dn}(\cdot;0)=1$ we get from (4.11) and (4.12) again $\lim_{k\to 1}a(k)=-\infty$ and $\lim_{k\to 0}a(k)=0$. Moreover, the mapping $k\in(0,1)\to g(k)=k\operatorname{sn}(L/c_{1};k)$ is strictly increasing (see Figure 13) with $k\operatorname{sn}(L/c_{1};k)>0$, $\lim_{k\to 1}k\operatorname{sn}(L/c_{1};k)=\tanh(\frac{L}{c_{1}})$ and $\lim_{k\to 0}k\operatorname{sn}(L/c_{1};k)=0$.
Once again, we split the analysis into subcases:
  1. (I)
    Let $Z>0$. Then we have:
    1. (i)
    2. (ii)
    This proves $(2)(i)$.
  2. (II)
  3. (III)
  1. (i)
  2. (ii)
  1. (i)
  2. ii)
The Proposition is now proved. ∎
@@
In the sequel we formally prove that, for the Neumann-Kirchhoff condition in (1.7), namely, $Z=0$, we get a single-lobe kink state for the sine-Gordon model with a profile on $[-L,L]$ given by formula in (4.8). For $Z\neq 0$ we found technical difficulties in determining the existence of these profiles (see Remark 4.10).
@@
Moreover, $a(k)>0$ is given univocally by relation in (4.11). For $Z<0$, there is not a single-lobe kink state with a first-component profile $\phi_{1,k}$ in (4.8) with $c_{2}$ satisfying $\frac{c_{1}}{2c_{2}}\geqq k_{0}$.

### v2

@@ 4.1 A priori restrictions on the parameter $Z$
$Z>\frac{\phi^{\prime}_{2}(L)}{\phi_{2}(L)}=-\frac{1}{c_{2}}\frac{y}{(1+y^{2})\arctan(y)}\equiv f(y),\;\;\quad y>0.$ (4.2)
@@
From (4.2) we get initially that $Z<\frac{1}{c_{2}}$ because $f(y)>-\frac{1}{c_{2}}$ for all $y>0$. Now, for $a\leqq 0$ we have $y=e^{-a/c_{2}}\geqq 1$. Then $Z>f(y)\geqq f(1)=-\frac{2}{\pi c_{2}}$, and so $Z<\frac{2}{\pi c_{2}}$. For $a>0$, we have $y<1$ and so $f(y)<f(1)$. Thus, for $a\to 0^{+}$, $f(y)\to f(1)$, and therefore $f(1)\leqq-Z$. But, for $a=0$, from (4.2) follows Z>f(1)-Z>f(1). In conclusion, we need to have the restriction $Z\in(-\infty,\frac{2}{\pi c_{2}})$. ∎
@@ 4.2 Existence of positive single-lobe states (subluminal librations)
$c_{1}^{2}\phi^{\prime\prime}_{1}(x)+\sin(\phi_{1}(x))=0,\qquad x\in[-L,L],$ (4.3)
with a single-lobe on [L,L][-L,L], namely, $\phi_{1}$ is even, the maximum of $\phi_{1}$ is achieved at a single internal point of [L,L][-L,L], namely, $x=0$, and $\phi^{\prime}(x)<0$ on $(0,L]$ (see Figure 3 and the phase-plane in Figure 5).
Figure 5: Phase-plane of a positive single-lobe kink state for (1.6). Here we take $c_{1}=c_{2}=1$ and the translation variable $\xi=x-ct$ with $c=0$ (subluminal waves). The picture shows the phase plane of the solutions $\phi=\phi(\xi)$ for (1.6). The separatrices are represented in red solid lines and the librations are the periodic waves inside the separatrices (in light gray). The single-lobe kink profile (depicted with blue thick arrow lines) is composed of a periodic libration, symmetric in [L,L][-L,L] and connecting $\phi=\pi$ with itself, glued with a decaying kink connecting $\phi=\pi$ to $\phi=0$ along a separatrix (color online).
@@
$\frac{c_{1}^{2}}{2}[\phi^{\prime}_{1}(x)]^{2}-\cos(\phi_{1}(x))=-\cos(\phi_{1}(0))\Longrightarrow-\frac{c_{1}^{2}}{2}[\phi^{\prime}_{1}(x)]^{2}+1-\cos(\phi_{1}(x))=E,\;\;x\in[-L,L],$ (4.4)
@@
In the sequel we obtain two explicit formulae for the profile $\phi_{1}$ in (4.7) being an even single-lobe profile on [L,L][-L,L] such that $\phi_{1}^{\prime}(x)<0$ for $x\in(0,L)$ and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. As we shall see, these formulae will depend on the position of $\phi_{2,a}(L)$ with regard to $\pi$, where the shift-parameter $a$ determines the kink-soliton profile in (1.8) (see Figure 3).
@@
Let $L$ and $c_{1}$ be fixed positive constants. We consider a priori $a<0$ in (1.8) which determines the kink-soliton profile $\phi_{2,a}$ (therefore $\phi_{2,a}(L)>\pi$). Suppose that $\phi_{1}$ satisfies equation in (4.3) and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. Then, there is a family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (4.3) of subluminal type with a single-lobe profile on [L,L][-L,L] given by,
@@
Next, from condition $\phi_{1}(x)>\pi$ on [L,L][-L,L], the property $\operatorname{sn}(2K(k);k)=0$ and (4.7), we need to have $\frac{x}{c_{1}}+K<2K$ for all $x\in[0,L]$ (see Figure 4 above). Therefore, a priori we get that modulus $k$ satisfies $K(k)>\frac{L}{c_{1}}$. Next, for $\frac{L}{c_{1}}\leqq\frac{\pi}{2}$ we obtain immediate a family of single-lobe solutions $k\to\phi_{1,k}$ for any $k\in(0,1)$. For $\frac{L}{c_{1}}>\frac{\pi}{2}$, we consider $k_{0}\in(0,1)$ such that $K(k_{0})=\frac{L}{c_{1}}$, then we have the family of single-lobe profiles $k\in(k_{0},1)\to\phi_{1,k}$ defined in (4.8). Lastly, from (4.3) we get immediately $\phi^{\prime\prime}_{1,k}(x)<0$ for all $x\in[-L,L]$. ∎
@@
We consider $L=\pi$, $c_{1}=1$, then $K(k)=\pi$ for $k\approx 0.9843\equiv k_{0}$, and so for $k>k_{0}$ we get the family of subluminal solutions $k\to\phi_{1,k}$ in (4.8) with a single-lobe profile on [L,L][-L,L]. Figure 4 shows the profile of $\phi_{1,k}$ for $k^{2}=0.98$.
@@
Let $L,c_{1}$ be fixed positive constants. We consider a priori $a>0$ in (1.8) which determines the kink-soliton profile $\phi_{2,a}$ (with $\phi_{2,a}(L)<\pi$). Suppose that $\phi_{1}$ satisfies equation in (4.3) and $\phi_{1}(-L)=\phi_{1}(L)=\phi_{2,a}(L)$. Then, there is a family of solutions $\phi_{1}\equiv\phi_{1,k}$ for equation in (4.3) of subluminal type with a single-lobe profile on [L,L][-L,L] given by
@@
From $a>0$, $\phi_{1}(0)>\pi$ and the boundary condition $\phi_{1}(L)=\phi_{2,a}(L)<\pi$, we have a unique $\in(0,L), b=b$(k)b=b(k), such that $\phi_{1}(b)=\pi$, and so for $x\in[0,b]$, $\phi_{1}(x)\in[\pi,\phi_{1}(0)]$, and for $x\in[b,L]$, $\phi_{1}(x)\in[\phi_{1}(L),\pi]$. Thus, from (4.7) we get the single-lobe profile in (4.9). Next, from $\phi_{1,k}(b)=\pi$ and the strictly-decreasing property of $\phi_{1,k}$ on $[0,L]$ we need to have $b(k)=c_{1}K(k)$ (see Figure 7 above). Therefore, $K(k)=\frac{b}{c_{1}}<\frac{L}{c_{1}}$. Moreover, from (4.3) we obtain $\phi^{\prime\prime}_{1}(b)=0$, $\phi^{\prime\prime}_{1}(x)<0$ for $x\in[0,b)$ and $\phi^{\prime\prime}_{1}(x)>0$ for $x\in[b,L]$ (see Figure 9 below). ∎
@@
We consider $L=\pi$, $c_{1}=1$, then $K(k)=\pi$ for $k\approx 0.984432\equiv k_{0}$, and so for $k\in(0,k_{0})$ we get the family of subluminal librational solutions $k\to\phi_{1,k}$ in (4.9) with a single-lobe profile on [L,L][-L,L] with $b=K(k)$. Figures 8 and 8 show the parts of $\phi_{1,k}$ in (4.9) when they are considered on the complete interval $0,\pi]$ with $k=0.5$. Figure 9 shows the full profile of $\phi_{1,k}$ in (4.9) with $k=0.5$, $b=K(0.5)$ and it which is based on gluing the two components for $\phi_{1,k}$ in (4.9) from the profiles in Figures 8 and 8. We note that the “corners” (namely, the points where the derivative does not exist) in Figures 8 and 8 occur simultaneously at $x=b=K(0.5)$.
@@
Figure 9: Graph of the single-lobe solution (in solid blue line), composed by the extension of $\phi_{1,k}$ in (4.9) with $c_{1}=1$, $k=0.5$ to the whole interval $x\in[-\pi,\pi]$. The glueing of both solutions occurs at $x=b$ and x=bx=-b with $b\approx 1.85407$ (color online).
@@ 4.3 Existence of a single-lobe kink state
Let $L,c_{1}$ fixed positive constants. For $a<0$ we consider the kink-soliton profile $\phi_{2,a}$ in (1.8) with $\phi_{2,a}(L)>\pi$ and the single-lobe profile $\phi_{1,k}$ in (4.8).
We have a single-lobe kink state, namely $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, in the following two cases:
@@
The proof of Proposition 4.7 shows formulae for $m_{0},m_{1},m_{2}$. Since $\tanh(K(k))>k$ for all $k\in(0,1)$, then in the case of single-lobe kink state existence $(1)(i)$ above we necessarily have $\frac{c_{1}}{2c_{2}}-k_{0}>0$.
As the proof of Proposition 4.7 is a bit technical, we will divide it into two lemmas, where the first of these shows that we can to glue the subluminal-profile $\phi_{1,k}$ in (4.8) with a kink-soliton profile $\phi=\phi_{2,a}$, $a=a(k)$, at the vertex $\nu=L$, namely, $\phi_{1,k}(L)=\phi_{2,a}(L)$ and with $k$ satisfying a priori $K(k)>\frac{L}{c_{1}}$. Here, we establish an explicit formula for the shift parameter $a=a(k)$. In the second Lemma, we determine the correct values of $k$ and $Z$ such that $(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, namely, such that we have the $\delta$-coupling relation $2\phi^{\prime}_{1,k}(L)=\phi^{\prime}_{2,a}(L)+Z\phi_{2,a}(L)$.
Lemma 4.9.
Let $L,c_{1},c_{2}$ fixed positive constants. For $a<0$ we consider the kink-soliton profile $\phi_{2,a}$ in (1.8) with $\phi_{2,a}(L)>\pi$ and the single-lobe profile $\phi_{1,k}$ in (4.8). Then, $\phi_{1,k}(L)=\phi_{2,a}(L)$ for $a=a(k)$ given by
where $k^{\prime}=\sqrt{1-k^{2}}$.
Proof.
Let $a<0$ and with $k$ satisfying a priori $K(k)>\frac{L}{c_{1}}$. By formulae (1.8), (4.8), and continuity condition $\phi_{1,k}(L)=\phi_{2,a}(L)$, we obtain the following first formula for the shift parameter $a=a(k)$,
@@
where $\mathrm{dn}(\cdot;k)$ and $\mathrm{cn}(\cdot;k)$ represent the Jacobian elliptic functions of dnoidal and cnoidal type, respectively (see Figure 7). Here, $k^{\prime}$ is defined by $k^{\prime}=\sqrt{1-k^{2}}$. In the following, we show formula in (4.11). Indeed, for $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$ we get from (4.7) that
$\cos(4\beta)=-1+2k^{2}sn^{2}\Big(\frac{L}{c_{1}}+K;k\Big).$
Then, from $\cos(4\beta)=2\cos^{2}(2\beta)-1$ follows $k^{2}\operatorname{sn}^{2}(\frac{L}{c_{1}}+K;k)=\cos^{2}(2\beta)$. Thus, since $2\beta\in[\frac{\pi}{2},\pi]$ and $\frac{L}{c_{1}}+K<2K$ we get $\cos(2\beta)=-k\operatorname{sn}(\frac{L}{c_{1}}+K;k)$ (see, for instance, Figure 4). Moreover, from $\cos^{2}(2\beta)+\sin^{2}(2\beta)=1$ and $1-k^{2}\operatorname{sn}^{2}=\mathrm{dn}^{2}$ (see [19]) we also obtain $\sin(2\beta)=\mathrm{dn}(\frac{L}{c_{1}}+K;k)$. Therefore,
@@
and so we get the property that the shift-value mapping $k\to a(k)$ will be smooth. Formula (4.13) shows that always we can glue a librational profile with a kink-soliton profile because $k^{\prime}\leqq\mathrm{dn}(x;k)$ for all $x$ and $k$. ∎
Lemma 4.10.
Let $L,c_{1},c_{2}$ fixed positive constants. For $a<0$ we consider the kink-soliton profile $\phi_{2,a}$ in (1.8) with $\phi_{2,a}(L)>\pi$ and the single-lobe profile $\phi_{1,k}$ in (4.8). Then, we have the $\delta$-coupling relation
$2\phi^{\prime}_{1}(L)=\phi^{\prime}_{2}(L)+Z\phi_{2}(L),$
with $\phi_{1}=\phi_{1,k}$, $\phi_{2}=\phi_{2,a(k)}$, and $a(k)$ satisfying relation in (4.13), provided that the conditions in items (1) and (2) of Proposition 4.7 hold. Moreover, there is not single-lobe kink state for the sine-Gordon model on a tadpole graph, namely $(\phi_{1,k},\phi_{2,a})\notin D_{Z}$, provided that the conditions in (3) and (4) of Proposition 4.7 hold.
The proof of Lemma 4.10 is a bit technical, so we leave this to be proven in Appendix.
@@
In the sequel we formally prove that, for the Neumann-Kirchhoff condition in (1.7), namely, $Z=0$, we get a single-lobe kink state for the sine-Gordon model with a profile on [L,L][-L,L] given by formula in (4.9). For $Z\neq 0$ we found technical difficulties in determining the existence of these profiles (see Remark 4.12).
@@
Moreover, $a(k)>0$ is defined uniquely by relation in (4.13). For $Z<0$, there is not a single-lobe kink state with a first-component profile $\phi_{1,k}$ in (4.9) with $c_{2}$ satisfying $\frac{c_{1}}{2c_{2}}\geqq k_{0}$.